Grade 11General Physics

A lift ascended with a constant acceleration of 4m/s square and then moved with a constant velocity and then retarted with 4m/s square. If the total distance asended is 20mts. and total time taken is 6 sec. then for how much time it travelled with constant velocity?

Priyansh Sharma
10 Years agoGrade 11

1 Answer

Charchit Tailong
10 Years ago
a=2m/sec2
v=4m/sec
d=20m
t=6sec

v=10(t)
v2-u2=2as
16=2*4*s.........so,s=8metres

remaining 12 metres
12=1/2(10t2)
root(2.4)